In my opinion, we ought to stop making our own drums and accept that outside suppliers offer, said Wim Niewindt, managingRequired: 1. Assuming that 95,000 drums are needed each year, what is the financial advantage (disadvantage) of buying the dr
Answer #1
Antilles refining
Part 1 - 95000 drums Make Buy Differential
Unit costs :-
Purchase Price 22 22
DM 10.6 -10.6
DL ( 8 * 80%) 6.4 -6.4
VOH ( 1.5 * 80%) 1.2 -1.2
Supervision 1 -1
Rent ( 285000 / 95000) 3 -3
Income / loss per unit 22.2 22 0.2
Finanicial advantage of buying the drums = 0.2 * 95000 = $19000
Part 2 - 100000 drums Make Buy Differential
Unit costs :-
Purchase Price 22 22
DM 10.6 -10.6
DL ( 8 * 80%) 6.4 -6.4
VOH ( 1.5 * 80%) 1.2 -1.2
Supervision 1 -1
Rent ( 285000 / 100000) 2.85 -2.85
Income / loss per unit 22.05 22 0.05
Finanicial advantage of buying the drums = 0.05 * 100000 = $5000
Part 3 - 125000 drums Make Buy Differential
Unit costs :-
Purchase Price 22 22
DM 10.6 -10.6
DL ( 8 * 80%) 6.4 -6.4
VOH ( 1.5 * 80%) 1.2 -1.2
Supervision 1 -1
Rent ( 285000 / 125000) 2.28 -2.28
Income / loss per unit 21.48 22 -0.52
Finanicial disadvantage of buying the drums = 0.52 * 125000 = $65000